# Toward Efficient Spiking Transformers: Synapse Pruning Meets Synergistic Learning-Based Compensation

### v2

@@ Methods
Method Architecture p [%] CR [%] Param [M] Accuracy [%]
LOCALZO+TET (Mukhoty et al. 2023) SEWResNet34 - - 63.47 78.58
IM-SNN (Hasssan et al. 2024) Resnet34 - - 21.27 74.42
IMP+TET (Shen et al. 2024) SEW-ResNet18 - - 63.47 78.70
EfficientLIF-Net (Kim et al. 2023) VGG16 - - 23.52 73.22
Spikformer (Zhou et al. 2023b) Spikformer-8-512-2048 - - 29.24/25.17 79.36
Table 1: Classification results on the ImageNet-100 dataset. The symbol denotes the baseline model. The terms p and CR indicate the pruning sparsity and compression ratio, respectively. The notation ‘/’ separates values measured on the entire model and on the ST blocks.
Dataset Method Architecture p [%] CR [%] Param [M] Accuracy [%]
CIFAR10
STBP-tdBN (Zheng et al. 2021) ResNet-19 - - 11.17 92.92
STP (Ma et al. 2025) ResNet18 - - 63.47 94.86
STL-SNN (Sun et al. 2023) ConvFC - - 11.37 92.42
Spikformer (Zhou et al. 2023b) Spikformer-4-384-1536 - - 9.32/7.08 95.19
CML (Zhou et al. 2023a) Spikformer-4-384-1536 - - 9.32/7.08 96.04
$\mathrm{L_{1}P}$+sLIF (ours) Spikformer-4-384-1536 80 60.62/79.94 3.67/1.42 93.94(-1.25)
$\mathrm{L_{1}P}$+sLIF (ours) Spikformer-4-384-1536 90 68.24/89.97 2.96/0.71 92.32(-2.87)
DSP+sLIF (ours) Spikformer-4-84-307 80 60.19/79.38 3.71/1.46 93.14(-2.05)
DSP+sLIF (ours) Spikformer-4-48-153 90 67.60/89.12 3.02/0.77 92.23(-2.96)
CIFAR10-DVS
TET (Deng et al. 2022) VGGSNN - - 9.54 77.33
STP (Ma et al. 2025) VGG11 - - 113.00 78.50
STL-SNN (Sun et al. 2023) ConvFC - - 1.53 77.30
Spikformer (Zhou et al. 2023b) Spikformer-2-256-1024 - - 2.59/1.58 80.90
CML (Zhou et al. 2023a) Spikformer-2-256-1024 - - 2.57/1.58 79.20
$\mathrm{L_{1}P}$+sLIF (ours) Spikformer-2-256-1024 80 48.65/79.75 1.33/0.32 78.00(-2.90)
$\mathrm{L_{1}P}$+sLIF (ours) Spikformer-2-256-1024 90 54.83/89.87 1.17/0.16 76.30(-4.60)
DSP+sLIF (ours) Spikformer-2-48-204 80 49.03/80.38 1.32/0.31 77.20(-3.70)
DSP+sLIF (ours) Spikformer-2-16-102 90 55.60/91.14 1.15/0.14 77.30(-3.60)
Table 2: Classification results on the CIFAR10 and CIFAR10-DVS datasets. Symbol represents the baseline model.
Experiments and Results
Experimental Settings
To evaluate the effectiveness of our method, experiments are conducted on both static and neuromorphic image datasets, ImageNet (Deng et al. 2009), CIFAR-10 (Krizhevsky, Hinton et al. 2009), and CIFAR10-DVS (Li et al. 2017). To further validate the generalization capability in downstream tasks, semantic segmentation is performed on the ADE20K dataset (Zhou et al. 2017). For a fair comparison, our method are directly applied to state-of-the-art (SOTA) pre-trained models implemented using the PyTorch and SpikingJelly frameworks (Zhou et al. 2023b; Yao et al. 2024; Fang et al. 2023). All experiments are executed on 4 NVIDIA A800 GPUs. The details can be seen in the Appendix.
Comparisons with SOTA Methods
Static Datasets Classification
We present our resulting lightweight models against other SNNs on the ImageNet-100 dataset (Tab. 1). The Spikformer-8-512-2048 model is selected as the baseline, which comprises 8 ST blocks with $d=512$ and $d_{\mathrm{m}}=2048$. Following pruning with a sparsity level of $p=90\%$, the number of model parameters is significantly reduced to below 77M, achieving a compression ratio (CR) exceeding $77\%$, while incurring only a minor drop in classification accuracy ($3.14\%$ for $\mathrm{L_{1}P}$+sLIF and $2.48\%$ for DSP+sLIF). When considering only the ST blocks, the CR exceeds $90\%$, with parameters reduced to below 2.52.5M. When the sparsity is substantially increased to an extreme value of $p=99\%$, the classification accuracy exhibits an approximate degradation of $16\%$. However, given that the parameter count of the ST blocks is reduced from 2424M to just 0.250.25M and 0.240.24M, the performance trade-off remains acceptable for lightweight deployment. Furthermore, compared to ResNet- or VGG-based convolutional models with comparable accuracy, our lightweight models still demonstrate significantly higher parameter efficiency.
On the CIFAR10 dataset (Tab. 2), ST-based models such as Spikformer and CML exhibit superior performance, with CML attaining the highest accuracy of $96.04\%$ when employing the Spikformer-4-384-1536 architecture. Utilizing our proposed $\mathrm{L_{1}P}$+sLIF method, an accuracy of $93.94\%$ is achieved under a sparsity level of $p=80\%$, accompanied by a substantial compression ratio of $60.62\%/79.94\%$, and a significantly reduced parameter count of 3.673.67M/1.421.42M. When the sparsity is further increased to $p=90\%$, the model maintains a competitive accuracy of $92.32\%$, while the parameter count is further reduced to 2.962.96M/0.710.71M. Similarly, the proposed DSP+sLIF strategy also demonstrates strong performance with enhanced parameter efficiency, achieving $93.14\%$ ($92.23\%$) accuracy with only 3.713.71M (3.023.02M) parameters under $p=80\%$ ($p=90\%$).
Neuromorphic Datasets Classification
On the CIFAR10-DVS dataset, the proposed lightweight models also demonstrate competitive performance, as summarized in Tab. 2. The baseline Spikformer model achieves an accuracy of $80.90\%$ with 2.592.59M/1.581.58M parameters. By applying the $\mathrm{L_{1}P}$+sLIF method under a sparsity level of $p=80\%$, the model is compressed to 1.331.33M/0.320.32M parameters while maintaining a commendable accuracy of $78.00\%$. Similarly, the DSP+sLIF strategy yields a compact model with the Spikformer-2-48-204 architecture, with only 1.321.32M/0.310.31M parameters and an accuracy of $77.20\%$ under the same sparsity level. The performance is similar while increasing the sparsity level.
Figure 4: Performance on the CIFAR10 and CIFAR10-DVS datasets under varying levels of pruning sparsity.
Figure 5: Convergence curves of pruned models on the CIFAR10 dataset during fine-tuning phase.
Ablation Analysis
Performance under different level of sparsity
To comprehensively evaluate the performance of the proposed lightweight strategies, we perform pruning experiments at varying pruning sparsity (denoted as ”sLIF”). For comparison, we also assess the accuracy of the original model after pruning (denoted as ”Pruned”). Additionally, to evaluate the contribution of synergistic learning, we conduct pruning experiments while retaining the original LIF neurons (denoted as ”LIF”). As illustrated in Fig. 4, model accuracy generally declines after pruning, with greater sparsity levels resulting in more pronounced performance degradation. Compared to the unstructured pruning strategy ($\mathrm{L_{1}P}$), the structured pruning approach (DSP) causes more substantial performance deterioration, which is consistent with prior findings. Nonetheless, fine-tuning is consistently an efficient strategy to recover performance. Notably, models fine-tuned with synergistic learning exhibit superior performance recovery, with the benefits of synergistic learning becoming increasingly evident at higher sparsity levels.
In addition to superior compensation performance, the convergence speed during the fine-tuning process is a critical factor in evaluating the effectiveness of a compensation strategy. We present the accuracy curves for the ‘sLIF’ and ‘LIF’ models under pruning sparsity of $p=0.99\%$ and $p=0.95\%$ on the CIFAR10 dataset, as shown in Fig. 5. In our experiments, we configured the fine-tuning process to run for 50 epochs. Across models pruned with various strategies, the sLIF approach consistently enables faster convergence during fine-tuning. Notably, this advantage is particularly pronounced for models pruned using the DSP method. Specifically, models fine-tuned with sLIF achieve convergence in approximately 20 epochs, whereas the ablation group employing LIF neurons requires around 40 epochs to reach convergence.
Impact on Model Inference Performance
Inference latency and energy consumption represent two fundamental metrics for assessing the inference performance of models. Given that structured pruning facilitates improved hardware accessibility, we evaluate the average batch inference runtime and theoretical power consumption of the DSP+LIF and DSP+sLIF models under pruning sparsity of $p=50\%$, $90\%$, and $99\%$, respectively. The relative ratios of the lightweight models compared to their corresponding baseline models are presented in Fig. 6.
SNN models predominantly utilize sparse accumulate operations as their primary computational units, resulting in significantly lower power consumption compared to traditional artificial neural networks. Additionally, the overall energy consumption of SNNs is influenced by both the model architecture and the neuronal firing rate. In this study, we estimate theoretical energy consumption following the method used in prior work (Yao et al. 2023b; Sun et al. 2024). As shown in Fig. 6, the relative energy consumption of pruned models, compared to the baseline, shows a positive correlation with the number of model parameters. Furthermore, sLIF models demonstrate slightly higher energy consumption than their LIF counterparts.
Figure 7: Representative examples of attention maps generated by the original (middle) and lightweight (right) models.
Visualization of Attention Maps
In the DSP+sLIF models, the dimensions of the $Q$, $K$, and $V$ representations are reduced to extremely low levels. For instance, pruning the original ’Spikformer-8-512-2048’ architecture (Tab. 1) with a pruning sparsity of $p=90\%$ results in a lightweight ’Spikformer-8-48-204’ model, where only 48 dimensions are used to represent the patch embeddings. Example attention maps at the final time step are shown in Fig. 7. Despite the reduced dimensionality, the lightweight model still effectively captures image regions relevant to classification tasks. These results demonstrate that our compensation strategy is an efficient and effective approach for compressing ST-based models. Detailed visualization method and further examples are provided in the Appendix.
Figure 8: Evolution of IP parameters of the $Q$ embedding. ’Start’ indicates the initial parameter values.
The Evolutionary Process of IP Parameters
To evaluate the contribution of IP parameters to model performance, we visualize the evolution of the $u_{\mathrm{th}}$ and $\tau$ associated with the $Q$ embedding in the first ST block on the CIFAR10 and CIFAR10-DVS datasets. As illustrated in Fig. 8, the magnitude of IP parameter adaptation is positively correlated with the sparsity level. Moreover, the evolution is more pronounced on the more challenging CIFAR10-DVS dataset. These results indicate that IP parameters play a more significant role in enhancing model performance under increased task difficulty. More results on the $K$ and $V$ embeddings can be found in the Appendix, which demonstrate similar conclusions.
[…]
Param [M] 29.71 22.05 17.12 7.00
Acc [%] 72.86 67.70 65.02 56.00
Table 3: Performance on the ImageNet-1k dataset.
Performance on the Large-Scale Dataset
We conduct experiments with the DSP+sLIF method on the more challenging ImageNet-1K dataset (Tab. 3). The model architecture and experimental settings are aligned with those used for ImageNet-100. Compared to ImageNet-100, a more significant performance drop is observed at high sparsity levels (e.g., $16.86\%$ on ImageNet-1K vs. $2.48\%$ on ImageNet-100 at $p=90\%$). This degradation is theoretically reasonable, as larger-scale datasets require larger model size according to scaling laws (Dosovitskiy et al. 2021). Nevertheless, the DSP+sLIF method remains effective, achieving competitive performance under extreme compression even on large-scale datasets.
Performance on the Downstream Tasks
To further demonstrate the effectiveness of the proposed DSP+sLIF method on complex downstream tasks, we apply it to semantic segmentation. The ST-based SDTv2 models (Yao et al. 2024) are used as the baseline, and the DSP+sLIF strategy is employed to compress the original models with a pruning sparsity of $p=90\%$. As shown in Tab. 4, the resulting lightweight ST-based semantic segmentation model achieves $26.69\%$ mIoU performance while requiring only 4.704.70M parameters. The qualitative visualization and discussion can be found in the Appendix.
Method ST Param mIoU
[M] [%]
MetaFormer (Yu et al. 2022) $\times$ 15.50 32.90
DeeplabV3 (Zhang et al. 2022b) $\times$ 68.10 42.10
SDTv2 (Yao et al. 2024) $\checkmark$ 9.42 30.14
DSP+sLIF (ours) $\checkmark$ 4.70 26.69
Table 4: Semantic segmentation on the ADE20K dataset.
Conclusion
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China (2023YFF1204200), in part by the STI 2030–Major Projects (2022ZD0208500), and in part by the Sichuan Science and Technology Program (2024NSFTD0032, 2024NSFJQ0004, and DQ202410).
Appendix
Appendix S1 Related Work
Efficient Spiking Transformer
ST-based models have garnered significant attention due to their potential for biologically plausible computation and energy-efficient processing of spatio-temporal data (Zhou et al. 2023b; Yao et al. 2023a). However, the high computational cost and memory overhead inherent in their complex dynamics necessitate efficiency-oriented enhancements. Prior work can be broadly categorized into structure-oriented and engineering-oriented approaches. Structure-Oriented Optimization: A primary direction for improving efficiency in ST blocks lies in architectural redesign. Inspired by model pruning strategies, researchers have explored eliminating redundant spatial and temporal components in attention mechanisms and feedforward layers, thereby reducing computational burden without substantially degrading accuracy (Zhou et al. 2024b; Kang et al. 2024). Another line of work proposes meta-architecture redesigns (Yao et al. 2024; Wang et al. 2025), which tailor model structures to the unique characteristics of spiking data. Furthermore, the integration of lightweight functional modules—such as linearized QK-value embedding and simplified gating mechanisms—has shown promise in improving representational efficiency with minimal performance trade-offs (Zhou et al. 2024a; Datta et al. 2025; Lu et al. 2025). Engineering-Oriented Solutions: In parallel, engineering-based methods addressing practical deployment issues have been proposed. Weight quantization, which reduces numerical precision, is a widely used technique to enable model deployment on edge devices. Researchers have incorporated quantization-aware training into ST-based models to lower the bit-width of weights and activations, thereby minimizing memory footprint and enabling deployment on low-power hardware (Qiu et al. 2025). Similarly, advances in specialized neuromorphic hardware—such as event-driven computing cores and custom-designed crossbar arrays—have significantly reduced latency and energy consumption for ST-based models (Kim et al. 2025). Despite their success, these methods tend to be model-dependent or require non-trivial system-level integration. This lack of universality limits their adaptability to diverse architectures and downstream tasks.
Bio-Inspired Spiking Neuron Models
The spiking neuron, as the fundamental computational unit, plays a critical role in determining the overall performance of SNNs (Zheng et al. 2024). Among various neuron models, the LIF model is one of the most widely adopted due to its simplicity and computational efficiency. By training LIF neurons with synaptic plasticity mechanisms (Wu et al. 2018; Davies et al. 2025), SNN models have achieved competitive performance on various tasks (Sun et al. 2024; Ding et al. 2022; Liu et al. 2021). Despite these advantages, the limited representational capacity of LIF in capturing complex spatio-temporal dynamics has motivated the development of more biologically inspired neuron models. In spiking neurons, intrinsic parameters (i.e., the membrane time constant, firing threshold, and resting potential) play a crucial role in shaping internal dynamics, including input integration, spike generation, and the refractory period (Zhang et al. 2022a). Based on the biological mechanisms underlying the firing threshold, numerous enhanced variants of the LIF model have been proposed, focusing on the threshold adaptation, dynamic thresholds, and subthreshold dynamics (Sun et al. 2023; Huang et al. 2024; Chen et al. 2022). Compared to the standard LIF neuron, these models offer a more biologically plausible depiction of threshold behavior, thereby enhancing the neuron’s ability to robustly represent spatial patterns. Furthermore, the membrane time constant governs the temporal integration of incoming signals (Pazderka 2024). By introducing heterogeneity in this parameter, recent studies have significantly improved the capacity of SNNs to capture spatial features across multiple time scales (Fang et al. 2021; Perez-Nieves et al. 2021; Zhang et al. 2025). A key challenge in enhancing LIF neuron models lies in the effective integration of the synergistic interactions between synaptic and intrinsic parameters. Prior studies have typically treated these parameters as independent components or optimized them using information-theoretic approaches, often overlooking their interdependence (Li and Li 2013). This not only limits the model expressiveness but also complicates the training process. Therefore, two critical questions remain open: how to incorporate synaptic and intrinsic plasticity in a unified and efficient framework, and how to enable synergistic learning between these two forms of plasticity.
Appendix S2 Experimental Details
This appendix provides detailed information about the experimental setup used to evaluate our method, including datasets, implementation frameworks, hardware specifications, and configuration details.
Datasets
  • ImageNet (Deng et al. 2009) is a large-scale static image dataset widely used for object recognition tasks, comprising approximately 1.28 million training images and 50,000 validation images across 1,000 classes. In our experiments, all images are resized to a resolution of $\times224$ pixels for consistency. To evaluate the classification performance of our method on high-resolution static images, we adopt the ImageNet-100 subset (Chun-Hsiao Yeh 2022), where the classes are selected following prior work. This subset serves as a benchmark for comparison with state-of-the-art (SOTA) models in conventional computer vision tasks. Additionally, the full ImageNet-1K dataset is employed to assess the scalability of our method under the scaling-law setting.
  • CIFAR-10 (Krizhevsky, Hinton et al. 2009) is a static image dataset consisting of 50,000 training images and 10,000 test images, each with a resolution of $\times32$ pixels and evenly distributed across 10 classes. It is used to evaluate the effectiveness of our method on low-resolution static images, offering a lightweight yet challenging benchmark for classification tasks.
  • CIFAR10-DVS (Li et al. 2017) is a neuromorphic dataset derived from the original CIFAR-10, recorded using a dynamic vision sensor (DVS). It comprises event streams from 10,000 samples spanning the same 10 classes as CIFAR-10, where each sample is represented as a sequence of address-event representations (AER) instead of conventional image frames. With temporal resolution on the order of microseconds, this dataset is well-suited for evaluating the performance of our method on event-based, neuromorphic vision tasks. In our experiments, each event stream is temporally averaged and segmented into 16 discrete time steps.
  • ADE20K (Zhou et al. 2017) is a widely used dataset for semantic segmentation, comprising 20,210 training images and 2,000 validation images annotated with 150 semantic categories. For segmentation tasks, images are typically resized to a resolution of $\times512$ pixels. Owing to its diverse scene compositions and fine-grained annotations, ADE20K serves as a benchmark to assess the generalization capability of our method on downstream tasks, particularly semantic segmentation.
Computing Infrastructure
All experiments are conducted on a machine with the following hardware and software configurations:
  • Hardware:
    • GPU: 4 $\times$ NVIDIA A800 80GB
    • CPU: Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz
    • Memory: 160 GB RAM
  • Software:
    • Operating System: CentOS Linux 7 (Core)
    • Python Version: Python 3.8.20
    • Deep Learning Frameworks: PyTorch 1.10.0+cu111 and SpikingJelly 0.0.0.0.12
The same environment is used for all experiments to ensure reproducibility.
Configuration Details
Tab. S5 summarizes the key configuration parameters used for each dataset in our experiments.

### v4

@@ III-B Synergistic Learning-based Compensation
Figure 4: Response re-alignment via synergistic plasticity of $u_{th}$ and $\tau$. Decreasing $u_{th}$ performs a horizontal shift of the f-I curve to the left, enabling the neuron to respond to weak inputs induced by pruning. Varying $\tau$ modulates the curvature and sensitivity of the response. The joint optimization of both parameters allows the sLIF neuron to maintain robust firing rates and information flow even under extreme pruning rates.
III-C Theoretical Analysis
III-C1 Synergistic Gradient Restoration
From an optimization perspective, synapse pruning induces a distribution mismatch between the membrane potential $u_{t}$ and the sensitive region of the surrogate gradient $\sigma^{\prime}(u_{t}-u_{\text{th}})$. The synergistic learning of both the firing threshold $u_{\text{th}}$ and the membrane time constant $\tau$ is necessary to effectively restore the gradient flow adaptively.
Let $\mathcal{S}=\{u\mid\sigma^{\prime}(u-u_{\text{th}})>\epsilon\}$ denote the gradient-sensitive window centered at $u_{\text{th}}$. Due to the reduced input current $\mathbb{E}[I_{t}]$ after pruning, the membrane potential $u_{t}$ drifts outside $\mathcal{S}$ (i.e., $u_{t}\ll u_{\text{th}}$), causing gradient vanishing. The restoration of gradients requires realigning the potential distribution with $\mathcal{S}$, which necessitates two complementary mechanisms. The synergistic learning of $\{u_{\text{th}},\tau\}$ functions as a dynamic affine transformation: $\tau$ amplifies the weak signals to a robust range (scaling), while $u_{\text{th}}$ fine-tunes the decision boundary (shifting). This coupling ensures that the membrane potential $u_{t}$ is optimally re-mapped into the gradient-sensitive window $\mathcal{S}$:
$\min_{u_{\text{th}},\tau}\sum_{t}\left|u_{t}(\tau)-u_{\text{th}}\right|\implies\sigma^{\prime}(u_{t}-u_{\text{th}})\gg 0.$ (22)
Thus, the joint optimization guarantees robust gradient recovery and efficient fine-tuning for pruned SNNs.
III-C2 Response Re-alignment via Synergistic Plasticity
From the perspective of neural dynamics, synapse pruning creates an “Input Distribution Shift” that impairs the information transmission capability of spiking neurons. We analyze this phenomenon using the frequency-current (f-I) curve and demonstrate how the synergistic plasticity of sLIF restores neuronal responsiveness. Based on the sLIF dynamics described in Eq. (11), the steady-state firing rate $f(I)$ for a constant input current $I$ can be derived as:
$f(I)=\left[t_{\text{ref}}+\tau\ln\left(\frac{I}{I-(u_{\text{th}}-u_{\text{rest}})}\right)\right]^{-1},$ (23)
where $t_{\text{ref}}$ is the refractory period. The term $I_{\text{rheo}}=u_{\text{th}}-u_{\text{rest}}$ represents the rheobase, defined as the minimum current required to trigger a spike. As shown in Fig. 4, synapse pruning significantly reduces the magnitude of the integrated current $I_{t}$, shifting its probability distribution towards zero. Consequently, the majority of inputs fall below the rheobase of a standard LIF neuron with a fixed threshold, resulting in a silent regime where $f(I)\to 0$ and the forward information flow is blocked.
The sLIF neuron recovers excitability by reshaping the f-I curve through the synergistic adaptation of $u_{\text{th}}$ and $\tau$. Specifically, learning a lower $u_{\text{th}}$ decreases $I_{\text{rheo}}$, which geometrically shifts the f-I curve to the left, ensuring that even weak pruned currents can exceed the firing threshold. Simultaneously, the membrane time constant $\tau$ acts as a scaling factor that modulates the curvature and gain of the response. While threshold adaptation ensures the neuron can fire, time constant adaptation regulates how it encodes signals. This synergistic coupling allows the sLIF neuron to dynamically match its dynamic range with the pruned input distribution, preserving robust information transmission in lightweight models.
TABLE I: Configuration Parameters for Experiments